OpenStudy (nessa.noodle):

How much of an alloy with 30% copper must be added to 25 pounds of a second alloy which is 20% copper to create a 27% alloy?

Hayhayz (hayhayz):

Welcome to OpenStudy! Please post your questions in the correct subject for others to help ☺

OpenStudy (sf_r):

(.15)(20) + (.10)(x) = (.12)(20 + x) 3 + .1x = 2.4 +.12x 3 - 2.4 = .12x - .1x .6 = .02x x = 30 pounds of Alloy 2 For the second problem, see if you can arrive at the key equation, (.30)(x) + (.20)(25) = (.27)(x + 25) Hope this helps.

ILovePuppiesLol (ilovepuppieslol):

please close this question now thanks ;)